# Accelerating Urban Science Research with AI Urban Scientist

### v1

@@ Abstract
Cities are complex, adaptive systems whose underlying principles remain difficult to disentangle despite unprecedented data abundance. Urban science therefore faces a fundamental challenge: converting vast, fragmented and interdisciplinary information into coherent explanations of how cities function and evolve. The emergence of AI scientists, i.e., agents capable of autonomous reasoning, hypothesis formation and data-driven experimentation, offers a new pathway toward accelerating this transformation, yet general-purpose systems fall short of the domain knowledge and methodological depth required for urban science research. Here we introduce a knowledge-driven AI Urban Scientist, built from hypotheses, peer-review signals, datasets and analytical patterns distilled from thousands of high-quality studies, and implemented as a coordinated multi-agent framework for end-to-end inquiry. The system generates structured hypotheses, retrieves and harmonizes heterogeneous datasets, conducts automated empirical analysis and simulation, and synthesizes insights in forms compatible with urban scientific reasoning. By providing reusable analytical tools and supporting community-driven extensions, the AI Urban Scientist lowers barriers to advanced urban analytics and acts not merely as an assistant but as an active collaborator in revealing the mechanisms that shape urban systems and in guiding the design of more resilient and equitable cities.
@@
Urban science plays a central role in understanding the dynamics of cities, diagnosing emerging urban challenges, and informing policy decisions on climate resilience, public health, economic development, and sustainability. Yet despite rapid growth in data availability—from remote sensing to mobility traces—and steady advances in statistical and computational methods, the research workflow in urban science remains highly manual and fragmented. Researchers typically begin with observation-driven ideas, search for relevant datasets using ad hoc heuristics, manually assemble heterogeneous data sources, write extensive custom code across multiple languages, and iterate on results with limited reproducibility. These challenges inhibit systematic discovery, constrain hypothesis generation, and slow the translation of urban data into actionable insights.
@@
Recent progress in large language models (LLMs) and autonomous agents has led to the emergence of the AI Scientist paradigm: systems capable of assisting or automating parts of the scientific process, including literature synthesis, hypothesis generation, dataset retrieval, code writing, and experiment execution. These systems promise to accelerate research by integrating reasoning, memory, and automation. However, general-purpose AI Scientist frameworks—largely developed for fields such as machine learning, chemistry, or physics—are not directly applicable to urban science. Urban phenomena are context-dependent, data are multi-source and heterogeneous, and analytical workflows require interdisciplinary knowledge spanning statistics, geospatial modeling, economics, epidemiology, and environmental science. Existing AI reviewer agents trained on ML corpora may misjudge high-quality urban science papers as “not novel enough,” reflecting a fundamental mismatch in epistemic standards. Similarly, existing data agents cannot extract datasets that only appear in scientific articles’ Data Availability sections, which represent some of the most important resources in urban research.
To address these gaps, we propose AI Urban Scientist (Figure 2), a domain-informed, multi-agent system designed specifically for urban science research. Our system incorporates the strengths of the AI Scientist paradigm while embedding domain-specific knowledge bases, workflow standards, and analytical conventions of urban research. The system is built around four core agents: 1) Ideation Agent, grounded in the CAMP framework (Context, Variables, Mechanism, Pattern), capable of recombining, transferring, or transforming hypothesis components extracted from 15K+ papers, and refining them through multi-disciplinary virtual scientists and a domain-trained critic. 2) Critic Agent, aligned with the standards of Nature and Nature Cities, fine-tuned on 2,000+ expert review reports and impact-factor–aligned hypothesis tiers to provide high-quality scientific judgement. 3) Data Search Agent, which constructs structured data cards by reading the Data Availability sections of 2,000+ urban science papers, integrates public and private datasets, and performs semantic similarity matching between hypotheses and datasets. 4) Data Analysis Agent, which uses a code base of over 10,000 scripts extracted from published studies to plan experiments, generate multi-language analysis pipelines, execute code, correct errors, and call simulators when empirical data are unavailable.
@@
Together, these agents create an end-to-end AI research workflow capable of generating hypotheses, discovering and integrating datasets, conducting empirical analysis, and producing interpretable scientific outputs. Beyond the autonomous workflow, we develop an open platform (Figure 8) that allows researchers to directly use individual agents, experiment with full pipelines, and contribute new tools, datasets, and analytical components. This community-driven ecosystem aims to establish shared standards for hypotheses, datasets, and analytical practices in urban science—accelerating research and improving reproducibility.
@@
  • We identify the potential of AI Scientists for urban science, highlight the limitations of general-purpose AI Scientist systems when applied to urban research, and articulate the need for domain-informed scientific agents.
  • We build AI Urban Scientist, a multi-agent system grounded in urban science knowledge bases and capable of performing the full research workflow.
  • We develop an open platform and ecosystem for community use and extension, enabling researchers to explore urban questions more systematically and collaboratively.
Traditional Urban Science Research
Contemporary Workflow and Challenges
Contemporary urban science research still relies heavily on a linear and highly manual workflow—researchers first generate observation-driven hypotheses, then manually search for suitable datasets, and finally hand-code analytical procedures. This traditional pipeline presents several structural limitations. Problem discovery remains narrow, as widely accepted urban science questions (e.g., scaling laws) dominate the field and constrain innovation. Hypothesis formation is often weak because it depends largely on expert experience and subjective observation rather than systematic exploration. Data remain underused due to the difficulty of locating relevant sources across heterogeneous domains, and analysis is inefficient because each study requires bespoke coding and processing. Moreover, reproducibility is low, as the field lacks the open-source tooling and standardized workflows common in machine learning. These challenges are amplified by the cross-disciplinary nature of urban science: integrating perspectives from geography, sociology, public health, and data science is difficult, and collaboration between domain experts and AI/CS researchers is often hindered by asymmetric knowledge gaps. Together, these barriers limit the ability of research teams to efficiently discover problems, generate robust hypotheses, and leverage the full potential of modern urban data.
Current Role of AI in Urban Science
Over the past two decades, the rise of large-scale digital traces—ranging from mobility records and social media to satellite imagery and IoT systems—has fundamentally shaped contemporary urban science, enabling a new data-intensive paradigm for understanding cities [1]. Despite this rapid growth, the actual role of AI in advancing core urban science research remains limited.
Most existing applications fall under urban computing, where machine learning models are applied to big-data streams to address urban management tasks such as traffic prediction, energy forecasting, or crowd-flow estimation [2, 3]. While effective in engineering contexts, these methods have not transformed how urban scientists generate hypotheses or develop theories. Their influence remains largely confined to technical prediction tasks rather than the broader scientific workflow of question discovery, hypothesis formation, and causal reasoning.
AI is also used when key urban indicators or first-hand measurements are difficult or costly to obtain—for instance, using machine learning to infer socioeconomic conditions from satellite images or to estimate mobility patterns from sparse data [4]. Such predictive models help fill data gaps but do not fundamentally shift conceptual frameworks or advance the theoretical core of urban science. More recently, large language models (LLMs) have begun to enter the research process. Researchers employ LLMs for literature exploration, preliminary methodological discussions, or experiment ideation [5]. However, these uses remain informal and are not yet systematically integrated into the scientific pipeline. Early visions of how LLMs could influence urban planning and computational urbanism are emerging [6], but these developments are still at an exploratory stage.
In summary, AI currently plays a supportive rather than transformative role in urban science. It accelerates certain tasks and fills data gaps but does not yet address deeper structural challenges such as limited problem discovery, weak hypothesis formation, inefficient analytical workflows, and persistent cross-disciplinary knowledge barriers. These limitations point to the need for a new research paradigm in which AI—potentially implemented as coordinated multi-agent systems—can collaborate with human researchers to explore data, formulate questions, test hypotheses, and integrate knowledge across disciplines, ultimately enabling a more systematic and discovery-driven mode of urban science.
@@
The Emergence of AI Scientist
The rapid advancement of large language models (LLMs) and autonomous AI agents has given rise to the concept of the AI Scientist—an AI system capable of assisting, accelerating, or even autonomously executing core components of the scientific workflow [7, 8]. Unlike earlier machine-learning applications focused mainly on prediction, AI Scientist systems integrate deep reasoning, automated coding, literature understanding, and iterative experimentation. Recent work demonstrates that LLM agents can generate research questions, synthesize theories, design experiments, and evaluate evidence [9, 10], while auto-coding capabilities translate ideas into executable analytical pipelines [11]. These advances have already begun reshaping fields such as data science [12], chemistry [13], and the social sciences [14]. Together, they suggest that AI Scientist systems could similarly transform urban science, enabling more automated, exploratory, and generative knowledge creation across its complex, data-rich research landscape.
How Can AI Scientist Help with Urban Science
Figure 1: The AI urban scientist workflow. An overview of how an AI Scientist system can assist urban research by automating each stage of the scientific pipeline—from identifying key topics, generating research hypotheses, discovering relevant datasets, and producing experimental analyses, to drafting the final research paper. This workflow illustrates the end-to-end integration of ideation, interdisciplinary knowledge synthesis, data retrieval, auto-coding, and result interpretation enabled by AI-driven research agents.
@@
The AI Scientist framework offers transformative opportunities for urban science by reshaping each step of the scientific workflow—from problem discovery to hypothesis building, data acquisition, analysis, and publication. As depicted in Figure 1, AI agents can operate as research collaborators capable of synthesizing knowledge, automating analysis, and bridging disciplinary divides [7, 8]. Below, we outline several key ways in which AI Scientist systems can advance urban science.
@@
  • Accelerating Ideation. Problem Discovery and Hypothesis Formation. Urban science has long suffered from a narrow band of canonical research problems and a heavy reliance on expert intuition. AI Scientist agents can explore large corpora of literature, extract underexplored topics, generate novel research directions, and synthesize interdisciplinary theories [9]. These systems can identify emerging urban challenges—ranging from environmental risks to socio-spatial inequalities—and propose testable hypotheses at a pace and breadth far beyond human capability. This capability is particularly valuable in urban settings, where questions often span geography, sociology, economics, environmental science, and engineering.
@@
  • Comprehensive and Automated Data Search. Identifying suitable datasets is often one of the most time-consuming steps in urban science. AI agents can perform broad data discovery across open data portals, scientific archives, satellite sources, and sensor datasets. Recent developments in LLM-guided dataset retrieval [12] allow the AI Scientist to match hypotheses with relevant data sources, assess data fitness, and even suggest complementary datasets to enrich analysis. This greatly reduces the manual burden and expands the scope of empirical studies.
  • Automated Experimental Design and Data Analysis. Through auto-coding and LLM-to-code translation capabilities [11], AI Scientist systems can automatically generate analytical pipelines—ranging from causal inference models and spatial econometrics to simulation workflows and machine learning pipelines for environmental prediction. This not only accelerates research execution but also increases reproducibility by producing self-documented code, structured outputs, and standardized evaluation procedures. In urban science—where data can be noisy, high-dimensional, or incomplete—AI Scientist systems can also propose alternative analyses or robustness checks autonomously.
  • Providing Reusable Tools and Accelerating Community Building. A major limitation of current urban science is the lack of shared computational tools and standardized workflows. AI Scientist agents can help create modular, reusable codebases, automated templates for data processing, and shared analytical components that reduce redundancy across research groups. This mirrors the open-source tradition of machine learning but brings it into the urban science domain, fostering community-building and improving reproducibility across studies [10].
In summary, AI Scientist systems offer a powerful means to overcome the structural challenges of contemporary urban science. By accelerating ideation, enabling interdisciplinary synthesis, enhancing data accessibility, automating analyses, and fostering tool-sharing, AI Scientist has the potential to transition the field from a manual, expert-driven paradigm toward a more systematic, scalable, and discovery-oriented mode of scientific inquiry.
@@
Figure 2: Our AI urban scientist system. The system integrates five core knowledge bases: (1) a hypothesis base distilled from 15k academic papers, (2) a review base containing 2k+ expert comments, (3) a data base of 20k+ urban datasets, (4) a code base of 10k+ analytical scripts, and (5) a simulator reflecting standard urban modeling practices. These components support four collaborating agents: an Ideation Agent that generates and mutates hypotheses, a Critic Agent fine-tuned on review knowledge, a Data Agent responsible for data discovery, and an Analyzing Agent that executes analytical workflows. Together, the system emulates the workflow of a domain-informed urban scientist, enabling reliable hypothesis generation, dataset matching, analysis execution, and scientific reasoning.
While recent progress in general-purpose AI scientist systems demonstrates impressive capabilities in hypothesis generation, literature synthesis, code writing, and autonomous experimentation, these systems remain insufficient for advancing the domain of urban science. The fundamental reason is that urban science differs from physics, machine learning, or chemistry in its theoretic development, data structures, methodological toolkit, and epistemic culture. Urban research requires extensive domain expertise to interpret heterogeneous datasets, understand contextual socioeconomic mechanisms, and apply appropriate analytical frameworks. Current AI Scientist agents lack these capabilities.
A striking example lies in state-of-the-art AI reviewers trained on machine learning peer-review corpora such as OpenReview. These agents can evaluate novelty and methodological rigor within ML domains, yet they fail when applied to interdisciplinary urban research. For instance, such models may incorrectly judge a Nature-level urban science paper as lacking novelty simply because it does not introduce a new algorithmic method—a misunderstanding rooted in their training on ML evaluation criteria, not on the epistemology of urban science. This mismatch illustrates that general-purpose AI reviewers and AI scientists are fundamentally misaligned with the intellectual standards and knowledge structures of urban research.
@@
This motivates the design of a knowledge-driven AI Urban Scientist, as illustrated in Figure 2. The system integrates a large hypothesis base distilled from 15K academic papers, a review base built from expert commentary, a curated data base of 20K+ urban datasets, and a code base and simulator representing standard analytical practices. Through coordinated agents—an Ideation Agent, Critic Agent, Data Agent, and Analyzing Agent—the system mirrors the workflow of a high-level human scientist: generating hypotheses, refining them through critique, identifying relevant datasets, executing analytical pipelines, and synthesizing insights. By embedding domain knowledge into each stage, the AI Urban Scientist is able to reason with the depth, specificity, and contextual awareness required for credible urban research.
Automatic Urban Science Ideation
@@
To generate high-quality and domain-grounded hypotheses for urban science, our system introduces the Ideation Agent, a knowledge-driven module designed around the CAMP theoretical framework. In urban science, a mature and scientifically testable hypothesis can typically be decomposed into five essential components:
@@
Through these structured transformations, the Ideation Agent produces an initial idea set. These ideas are then passed into a debate stage involving multi-disciplinary Virtual Scientists, whose diverse perspectives help refine the conceptual plausibility. In parallel, the Critic Agent, fine-tuned on urban science review corpora, evaluates methodological soundness and theoretical rigor. Together, these iterative feedback loops support a dynamic hypothesis evolution process, converging on refined, domain-consistent hypotheses suitable for empirical testing.
Reliable Urban Science Critic
To ensure that the hypotheses generated by the Ideation Agent are evaluated through a lens consistent with the standards of high-quality urban science, we develop a Critic Agent that emulates the judgment patterns of expert reviewers in the field. Unlike general-purpose AI reviewers, which tend to inherit evaluation criteria from machine learning venues such as OpenReview, our Critic Agent is explicitly grounded in the epistemic standards of Nature and Nature Cities, two flagship journals representing the highest level of urban and interdisciplinary research. As shown in Figure 3, the key steps of constructing the Critic Agent are as follows.
System-Level Alignment with Urban Science Standards.
We begin by constructing a domain-specific system prompt that integrates the editorial requirements, aims and scope, and mission statements of Nature and Nature Cities. These statements emphasize conceptual significance, interdisciplinary depth, empirical grounding, methodological rigor, and societal relevance. By embedding these principles directly into the system prompt, we ensure that the Critic Agent evaluates hypotheses according to norms that are aligned with the expectations of top-tier urban science journals.
Fine-Tuning with Expert Review Corpora.
Impact-Factor–Aligned Evaluation Calibration.
Through system prompt alignment, expert-review fine-tuning, and impact-factor–based calibration, the Critic Agent develops evaluation capabilities that reflect the values of the urban science community. It can judge whether a hypothesis has conceptual significance, cross-disciplinary relevance, empirical feasibility, and theoretical credibility. Its output includes both a tier rating (Tier 1–4 or reject) and explanatory feedback, enabling the hypothesis refinement loop to converge toward scientifically meaningful ideas.
Figure 4: The construction and training pipeline of the Critic Agent. We align the Critic Agent with domain standards by integrating the editorial principles of Nature and Nature Cities as the system prompt. A corpus of 15K papers provides 15K+ seed hypotheses, while 2K+ expert reviewer comments from Nature family journals supply high-quality positive and negative evaluation signals for fine-tuning. Additionally, hypotheses extracted from the 15K papers are labeled into four tiers according to journal impact factor, enabling multi-level calibration. Together, these components produce a domain-informed Critic Agent capable of assigning idea-quality tiers (Tier 1–4 or reject) consistent with expert judgment in urban science.
@@
Figure 5: The dataset construction and retrieval workflow of the Data Agent. The Data Agent extracts dataset information directly from the Data Availability sections of urban science papers using LLM-based semantic parsing, producing standardized data cards that include dataset name, region, time period, description, and URL. These cards populate a unified data pool spanning four major urban data categories: statistical infrastructure, human behavior, policy and survey, and multimodal sensing. Through semantic similarity matching, the Data Agent links hypotheses to relevant datasets and enables automated downloading, preprocessing, and integration for downstream analysis.
High quality empirical urban science depends on diverse data sources that cover infrastructure, mobility, environment, economy, health, policy, and human behaviour. However, many high value datasets are not directly discoverable through standard web search. They are often only mentioned in the data availability sections of academic articles, embedded in supplementary materials, or hosted on domain specific repositories and institutional portals. To address this gap, our AI urban scientist includes a Data Agent (Figure 5) that is responsible for discovering, structuring, and retrieving urban datasets in a unified way.
@@
The Data Agent first scans scientific articles, with a focus on their data availability statements and related sections. Using the semantic retrieval and information extraction capabilities of large language models, it identifies text segments that describe datasets, including dataset names, spatial coverage, temporal range, measurement types, and access links. Based on more than 2,000 domain papers, the agent converts these unstructured descriptions into structured data cards. Each data card contains
  • Name: Dataset name or short title;
  • Country / region: Main spatial coverage;
  • Time: Observation period or reference years;
  • Type: Data category and subcategory;
  • Description: Brief summary of variables, granularity, and use cases;
  • URL: Link to repository, DOI landing page, or access portal.
These data cards form a machine readable index that populates a unified data pool.

### v2

# AI Urban Scientist: Multi-Agent Collaborative Automation for Urban Research

@@ Abstract
Urban research aims to understand how cities operate and evolve as complex adaptive systems. With the rapid growth of urban data and analytical methodologies, the central challenge of the field has shifted from data availability to the integration of heterogeneous data into coherent, verifiable urban knowledge through multidisciplinary approaches. Recent advances in AI, particularly the emergence of large language models (LLMs), have enabled the development of AI scientists capable of autonomous reasoning, hypothesis generation, and data-driven experimentation, demonstrating substantial potential for autonomous urban research. However, most general-purpose AI systems remain misaligned with the domain-specific knowledge, methodological conventions, and inferential standards required in urban science. Here, we introduce the AI Urban Scientist, a knowledge-driven multi-agent framework designed to support autonomous urban research. Grounded in hypotheses, peer-review feedback, datasets, and research methodologies distilled from large-scale prior studies, the system constructs structured domain knowledge that guides LLM-based agents to automatically generate hypotheses, identify and integrate multi-source urban datasets, conduct empirical analyses and simulations, and iteratively refine analytical methods. Through this process, the framework synthesizes new insights in urban science and accelerates the urban research lifecycle.
@@
More than half of the global population live in cities and therefore it is crucial to understand the dynamics of cities and diagnose emerging urban challenges, informing policy decisions on climate resilience, public health, economic development, and sustainability [1, 2]. Yet despite rapid growth in data availability from remote sensing to mobility traces, and steady advances in statistical and computational methods, urban research workflow remains highly manual and fragmented. Researchers typically begin with observation-driven ideas, search for relevant datasets using ad hoc heuristics, manually assemble heterogeneous data sources, write extensive custom code across multiple languages, and iterate on results with limited reproducibility. These challenges inhibit systematic discovery, constrain hypothesis generation, and slow the translation of urban data into actionable insights.
Recent progress in large language models (LLMs) and autonomous agents has led to the emergence of the AI Scientist paradigm [3, 4]: An AI scientist refers to an artificial intelligence–based system designed to conduct scientific research by executing the core steps of the scientific method with minimal human intervention. Its objective is to autonomously perform the iterative cycle of hypothesis formulation, solution or experimental design, experiment execution, result evaluation, and adaptive refinement, and in some cases to generate complete scientific manuscripts. Through feedback-driven learning, an AI scientist continuously improves its reasoning and experimental strategies, analogous to how human scientists iteratively test and refine hypotheses. In this sense, AI scientists represent a conceptual shift in artificial intelligence from computational tools that assist scientific tasks toward systems capable of autonomous scientific discovery.
@@
Such systems also hold the potential to accelerate urban research by integrating search, reasoning, coding, and related capabilities. However, general-purpose AI Scientist frameworks—largely developed for fields such as machine learning, chemistry, or physics—are not directly applicable to urban studies. Urban phenomena are context-dependent, data are multi-source and heterogeneous, and analytical workflows require interdisciplinary knowledge spanning statistics, geospatial modeling, economics, epidemiology, and environmental science. Existing AI reviewer agents trained on ML corpora may misjudge high-quality urban study papers as “not novel enough,” reflecting a fundamental mismatch in epistemic standards. Similarly, existing data agents cannot extract datasets that are mainly introduced by scientific articles, which represent some of the most important resources in urban research.
To address these gaps, we propose AI Urban Scientist (Figure 1), a domain-informed, multi-agent system designed specifically for urban research. At its core, the AI Urban Scientist is designed as a knowledge-driven, multi-agent system in which specialized agents collaborate to emulate the full scientific cycle in urban research. Instead of relying on a monolithic LLM, the system decomposes the research process into interacting agents responsible for hypothesis generation, data discovery, empirical analysis, result synthesis, and critical evaluation. These agents are coordinated through structured domain knowledge distilled from prior literature, peer-review practices, urban datasets, and standard analytical workflows, enabling iterative feedback and role-specific reasoning. By explicitly modeling scientific research as a collaborative and self-correcting multi-agent process aligned with established urban study norms, the AI Urban Scientist supports autonomous yet verifiable urban science and advances scalable and insightful urban research.
@@
Beyond the autonomous workflow, we develop an open platform (Figure 7) that allows researchers to directly use individual agents, experiment with full pipelines, and contribute new tools, datasets, and analytical components. This community-driven ecosystem aims to establish shared standards for hypotheses, datasets, and analytical practices in urban science—accelerating research and improving reproducibility.
Figure 1: Overview of the AI Urban Scientist system. The system supports autonomous urban research by automating each stage of the scientific workflow, including identifying key topics, generating research hypotheses, discovering and integrating relevant datasets, conducting empirical analyses and simulations, and drafting the final research manuscript. To enable this process, the system is grounded in four domain-specific knowledge bases: (1) a hypothesis base distilled from over 15,000 academic papers, (2) a review base comprising more than 2,000 expert peer-review comments, (3) a data base containing over 60,000 urban datasets, and (4) a code base of more than 10,000 analytical scripts and simulators reflecting standard urban modeling practices. These knowledge resources coordinate five collaborative agents: an Ideation Agent for hypothesis generation and mutation, a Search Agent for multi-source data discovery, an Analysis Agent for executing analytical workflows, a Writing Agent for synthesizing results into a research paper, and a Critic Agent, fine-tuned on review knowledge, for evaluating research quality and providing iterative feedback.
@@
  • We present the first AI Urban Scientist system that enables the entire urban research lifecycle—from problem formulation and hypothesis generation to data discovery, empirical analysis, and manuscript drafting—through autonomous and collaborative multi-agent coordination.
  • We propose a knowledge-driven multi-agent architecture grounded in structured urban science knowledge bases, including prior literature, peer-review practices, datasets, and analytical workflows, which systematically enhances domain alignment, methodological rigor, and research reliability.
  • We develop an open and extensible research platform that supports community use and further extension, enabling urban researchers to explore scientific questions in a more systematic, scalable, and collaborative manner.
@@
What Is an AI Scientist and Why Does It Matter?
An AI Scientist refers to an AI system designed to autonomously participate in the scientific discovery process by performing core research functions that traditionally rely on human expertise. Unlike conventional AI tools that support isolated tasks such as prediction or data analysis, an AI Scientist operates across the full scientific workflow, including literature synthesis, hypothesis generation, experimental design, data-driven reasoning, and iterative refinement of scientific knowledge.
The defining advantage of an AI Scientist lies in its ability to overcome fundamental human bottlenecks in scale, cognition, and coordination:
First, AI Scientist systems can systematically process and synthesize vast bodies of scientific literature at a speed and scope far beyond human capacity. By scanning tens of thousands of publications across multiple disciplines, these systems can identify latent patterns, unresolved contradictions, and underexplored research directions, enabling the generation of large numbers of testable hypotheses—including non-intuitive hypotheses that may be overlooked due to human cognitive biases or disciplinary conventions.
Second, AI Scientist systems excel at navigating high-dimensional and complex research spaces. In scientific settings characterized by large parameter spaces and interacting variables, human intuition often proves inefficient or unreliable. Through methods such as Bayesian optimization, reinforcement learning, and large-scale simulation, AI Scientist systems can autonomously design, optimize, and adapt experimental strategies, efficiently exploring vast search spaces that would otherwise be intractable.
Third, AI Scientist systems are inherently suited for large-scale pattern recognition and model construction in complex data environments. Modern scientific datasets are increasingly high-dimensional, noisy, and heterogeneous, posing significant challenges to human reasoning. AI systems can uncover nonlinear relationships and deep structural regularities across diverse data modalities, automatically constructing predictive and explanatory models that extend beyond human perceptual limits.
A further distinguishing feature of AI Scientist systems is their capacity for massively parallel and collaborative intelligence. Rather than functioning as a single monolithic agent, an AI Scientist can instantiate large populations of interacting agents with diverse reasoning strategies, enabling parallel exploration of hypotheses, analytical approaches, and interpretations. This form of scalable collective intelligence far exceeds the coordination capacity of human research teams.
In addition, AI Scientist systems are not constrained by social, institutional, or reputational pressures that shape human scientific behavior. Free from academic hierarchies, prevailing paradigms, and funding incentives, these systems can explore unconventional or counterintuitive research paths without psychological or professional cost, potentially opening avenues for disruptive discovery.
Finally, AI Scientist systems possess a unique capacity for continuous learning and deep interdisciplinary integration. Their knowledge bases can be updated in real time as new scientific results emerge, enabling cumulative and non-depreciating learning at a global scale. Moreover, by integrating knowledge across traditionally siloed disciplines, AI Scientist systems can reason from genuinely interdisciplinary perspectives, supporting scientific inquiry that transcends individual domain expertise.
Taken together, these capabilities define the AI Scientist not merely as a faster analytical tool, but as a new class of scientific agent—one that augments human intelligence by operating at scales, depths, and levels of integration that are fundamentally unattainable by human researchers alone.
How Can AI Scientist Help with Urban Research
@@
The AI Scientist framework offers transformative opportunities for urban research by reshaping each step of the scientific workflow—from problem discovery to hypothesis building, data acquisition, analysis, and publication. As depicted in Figure 1, AI agents can operate as research collaborators capable of synthesizing knowledge, automating analysis, and bridging disciplinary divides [5, 6]. Below, we outline several key ways in which AI Scientist systems can advance urban research.
@@
  • Accelerating Ideation. Problem Discovery and Hypothesis Formation. urban research has long suffered from a narrow band of canonical research problems and a heavy reliance on expert intuition. AI Scientist agents can explore large corpora of literature, extract underexplored topics, generate novel research directions, and synthesize interdisciplinary theories [7]. These systems can identify emerging urban challenges—ranging from environmental risks to socio-spatial inequalities—and propose testable hypotheses at a pace and breadth far beyond human capability. This capability is particularly valuable in urban settings, where questions often span geography, sociology, economics, environmental science, and engineering.
@@
  • Comprehensive and Automated Data Search. Identifying suitable datasets is often one of the most time-consuming steps in urban research. AI agents can perform broad data discovery across open data portals, scientific archives, satellite sources, and sensor datasets. Recent developments in LLM-guided dataset retrieval [9] allow the AI Scientist to match hypotheses with relevant data sources, assess data fitness, and even suggest complementary datasets to enrich analysis. This greatly reduces the manual burden and expands the scope of empirical studies.
  • Automated Experimental Design and Data Analysis. Through auto-coding and LLM-to-code translation capabilities [10], AI Scientist systems can automatically generate analytical pipelines—ranging from causal inference models and spatial econometrics to simulation workflows and machine learning pipelines for environmental prediction. This not only accelerates research execution but also increases reproducibility by producing self-documented code, structured outputs, and standardized evaluation procedures. In urban research—where data can be noisy, high-dimensional, or incomplete—AI Scientist systems can also propose alternative analyses or robustness checks autonomously.
  • End-to-end research workflow with closed-loop improvement. Beyond autonomously generating research ideas and conducting data-driven analyses, the AI Scientist can produce structured scientific feedback that explicitly identifies weaknesses in research design and analytical strategies. This feedback is then leveraged to iteratively refine hypotheses, experimental setups, and analytical pipelines, forming a closed-loop process of continuous improvement across the entire research workflow.
  • Providing Reusable Tools and Accelerating Community Building. A major limitation of current urban research is the lack of shared computational tools and standardized workflows. AI Scientist agents can help create modular, reusable codebases, automated templates for data processing, and shared analytical components that reduce redundancy across research groups. This mirrors the open-source tradition of machine learning but brings it into the urban research domain, fostering community-building and improving reproducibility across studies [11].
In summary, AI Scientist systems offer a powerful means to overcome the structural challenges of contemporary urban research. By accelerating ideation, enabling interdisciplinary synthesis, enhancing data accessibility, automating analyses, and fostering tool-sharing, AI Scientist has the potential to transition the field from a manual, expert-driven paradigm toward a more systematic, scalable, and discovery-oriented mode of scientific inquiry.
@@
While recent progress in general-purpose AI scientist systems demonstrates impressive capabilities in hypothesis generation, literature synthesis, code writing, and autonomous experimentation, these systems remain insufficient for advancing the domain of urban research. The fundamental reason is that urban research differs from physics, machine learning, or chemistry in its theoretic development, data structures, methodological toolkit, and epistemic culture. Urban research requires extensive domain expertise to interpret heterogeneous datasets, understand contextual socioeconomic mechanisms, and apply appropriate analytical frameworks. Current AI Scientist agents lack these capabilities.
A striking example lies in state-of-the-art AI reviewers trained on machine learning peer-review corpora such as OpenReview. These agents can evaluate novelty and methodological rigor within ML domains, yet they fail when applied to interdisciplinary urban research. For instance, such models may incorrectly judge a Nature-level urban research paper as lacking novelty simply because it does not introduce a new algorithmic method—a misunderstanding rooted in their training on ML evaluation criteria, not on the epistemology of urban research. This mismatch illustrates that general-purpose AI reviewers and AI scientists are fundamentally misaligned with the intellectual standards and knowledge structures of urban research.
@@
This motivates the design of a knowledge-driven AI Urban Scientist, as illustrated in Figure LABEL:fig:2. The system integrates a large hypothesis base distilled from 15K academic papers, a review base built from expert commentary, a curated data base of 20K+ urban datasets, and a code base and simulator representing standard analytical practices. Through coordinated agents—an Ideation Agent, Critic Agent, Data Agent, and Analyzing Agent—the system mirrors the workflow of a high-level human scientist: generating hypotheses, refining them through critique, identifying relevant datasets, executing analytical pipelines, and synthesizing insights. By embedding domain knowledge into each stage, the AI Urban Scientist is able to think with the depth, specificity, and contextual awareness required for credible urban research.
Automatic Urban Research Ideation
@@
To generate high-quality and domain-grounded hypotheses for urban research, our system introduces the Ideation Agent, a knowledge-driven module designed around the CAMP theoretical framework. In urban research, a mature and scientifically testable hypothesis can typically be decomposed into five essential components:
@@
Through these structured transformations, the Ideation Agent produces an initial idea set. These ideas are then passed into a debate stage involving multi-disciplinary Virtual Scientists, whose diverse perspectives help refine the conceptual plausibility. In parallel, the Critic Agent, fine-tuned on urban research review corpora, evaluates methodological soundness and theoretical rigor. Together, these iterative feedback loops support a dynamic hypothesis evolution process, converging on refined, domain-consistent hypotheses suitable for empirical testing.
@@
Figure 3: The dataset construction and retrieval workflow of the Data Agent. The Data Agent extracts dataset information directly from the Data Availability sections of urban research papers using LLM-based semantic parsing, producing standardized data cards that include dataset name, region, time period, description, and URL. These cards populate a unified data pool spanning four major urban data categories: statistical infrastructure, human behavior, policy and survey, and multimodal sensing. Through semantic similarity matching, the Data Agent links hypotheses to relevant datasets and enables automated downloading, preprocessing, and integration for downstream analysis.
High quality empirical urban research depends on diverse data sources that cover infrastructure, mobility, environment, economy, health, policy, and human behaviour. However, many high value datasets are not directly discoverable through standard web search. They are often only mentioned in the data availability sections of academic articles, embedded in supplementary materials, or hosted on domain specific repositories and institutional portals. To address this gap, our AI urban scientist includes a Data Agent (Figure 3) that is responsible for discovering, structuring, and retrieving urban datasets in a unified way.
@@
The Data Agent first scans scientific articles, with a focus on their data availability statements and related sections. Using the semantic retrieval and information extraction capabilities of large language models, it identifies text segments that describe datasets, including dataset names, spatial coverage, temporal range, measurement types, and access links. Based on more than 2,000 domain papers, the agent converts these unstructured descriptions into structured data cards.
Table 1: Schema of structured data cards used to index urban datasets.
Field Description
Name Dataset name or short descriptive title.
Country / Region Primary spatial coverage of the dataset.
Time Observation period or reference years covered by the data.
Type Data category and subcategory (e.g., environmental, socioeconomic, mobility).
Description Brief summary of key variables, spatial and temporal granularity, and typical use cases.
URL Link to the data repository, DOI landing page, or official access portal.
As show in Table 1, these data cards form a machine readable index that populates a unified data pool.